# Compact Cauchy horizons admit constant surface gravity

### v1

@@ 2 The cohomological equation
The foliated cohomology can be defined for more general foliated manifolds but is particularly simple to describe in this flow setting [11]. Let $\chi$ be a 1-form whose kernel is supplementary at each point to the tangent space $D:=\textrm{span}(n)$ of the leaves and $\chi(n)=1$. Then
@@
The literature on the cohomological equation is extensive but results are focused on proving the existence of $f$ for any $\kappa$ up to some normalization conditions to be imposed on the integral of $\kappa$. In general, the problem can be solved under restrictive conditions on the flow, in fact Katok’s conjecture (now proved for $H$ of dimension 3 [34, 28, 16], which is the physical 4-dimensional spacetime case) states that on a connected orientable compact manifold $H$ the equation $n(f)=\kappa-c(\kappa)$ can be solved for any function $\kappa$ and for some constant $c(\kappa)$, only if the flow is conjugate to a Diophantine linear flow on the torus.
Instead, in our spacetime problem we need to solve the cohomological equation for just one function $\kappa$. We do not expect severe constraints on the topology of $H$ or on the flow but we have two simplifications on our side. On the one hand, we already know the flow to be Riemannian and, on the other hand, the function $\kappa$ is not any function as it reads $\kappa=\omega(n)$ i.e. it is derived from a 1-form $\omega$ with special properties. We shall need to use these facts. In particular, we shall need [19, Lemma 5]
Proposition 2.2.
Assume the dominant energy condition. Then on the horizon $H$
${\rm d}\omega(n,\cdot)=0.$
As we shall see, previous studies on the cohomological equation [11] will really be used just in dealing with the exceptional non-isometric case $T^{3}_{A}$ (see case (iii-b) of Thm. 1.4) as it involves a Diophantine linear flow on the torus.
@@ 3 Consequences of ergodic theory and ribbon argument
Proof.
The flow $(\phi_{t})_{t\in\mathbb{R}}$ of $n$ is a measure preserving flow since it preserves the Riemannian volume element, in fact, as shown in the proof of [23, Prop. 22] $\widetilde{\textrm{div}}(n)=0$. Let us denote $\kappa=\omega(n)$. Then, Birkhoff’s Ergodic Theorem for measure preserving transformations gives [6, 42, 3]
@@
It can be observed that the argument of the limit is bounded as $\omega(n)$ is a continuous function on $H$, hence bounded. The important point is that by Birkhoff’s theorem this limit converges for some point. Let $p$ be one such point and suppose that the limit is zero. Let $n^{*}$ be a smooth 1-form field such that $n^{*}(n)=1$. A curve with tangent in $\ker n^{*}$ is said to be $n^{*}$-horizontal. The limit $\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\omega(\dot{x}){\rm d}t$ does not really depend on the starting point of the curve, so for any sufficiently close integral curve $x^{\prime}$ we can redefine the starting point $x^{\prime}(0)$ of the parametrization so that there is a horizontal curve $\sigma_{0}$ connecting $x(0)$ to $x^{\prime}(0)$ (this can be done so that its projection on a local cylinder neighborhood is a geodesic for $g_{T}$, see [19]). Now, it is possible to construct a ribbon from the curves $x|_{[0,t]}$, $\sigma_{0}|_{[0,1]}$ and $x^{\prime}|_{[0,t^{\prime}(t)]}$, and application of Stokes’s theorem and the bound described in [19] implies that
@@
The bound on the right-hand side does not depend on how elongated is the ribbon (see [19, Eq. (18)-(19)] and take into account that this conclusion only depends on $L_{n}g_{T}=0$ not on the spacetime interpretation of the various objects in that work) and in [19, App.] it is detailed how $t\to+\infty$ implies $t^{\prime}\to\infty$, and conversely. In fact, on the same appendix (footnote 6) it is shown that there is a constant $C>0$ such that $te^{-C}\leq t^{\prime}\leq te^{C}$. Dividing the previous equation by $t$ and taking the limit we get, as $t^{\prime}/t$ remains bounded away from zero, $\lim_{t^{\prime}\to\infty}\frac{1}{t^{\prime}}\int_{x^{\prime}([0,t^{\prime}])}\omega=0$. This shows that the set $O$ of points $p$ for which the orbit starting from it has first integral existing and equal to zero is open. In fact it is also closed, for if $q\in\partial O$ then for every neighborhood $U\ni q$, we can find some $p\in O$ for which the integral exists and vanishes. Then, reapplying the ribbon argument, we get that the same is true for the orbit from $q$. By the connectedness of $H$ we conclude $O=H$. By Birkhoff’s theorem $\int_{H}\omega(n){\rm d}\tilde{g}=0$. This shows that (a) implies its strengthened form and (c).
@@
Let us prove (a) implies (d). In the proof of [23, Prop. 24] we proved that for $N=H$ or $N$ equal to a torus closure of an orbit, the restricted flow of $n$ is volume preserving. Actually, the proof applies unaltered for any saturated submanifold $N$. The Birkhoff’s Ergodic Theorem can thus be applied to the submanifold which proves (d). Conversely, when the last equation of point (d) is applied to $N=H$ we get point (c), thus (d) implies (c). ∎
Remark 3.2.
@@
Theorem 3.3.
Suppose that the spacetime satisfies the dominant energy condition and let $n$ be any smooth lightlike field tangent to a compact degenerate horizon $H$. Let $\omega$ be the 1-form defined by Eq. (1). Then interpreting $H$ as a Riemannian flow structure, the properties of Theorem 3.1 are satisfied.
Proof.
The limit $\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\kappa(x(t)){\rm d}t$ is known to exist for some generator but it cannot be different from zero. Indeed, if it is negative then for every generator $\int_{0}^{t}\kappa(x(t)){\rm d}t=-\infty$ and by the equivalence proved in [19, Def. 4] the generators of the horizons would be all future incomplete. Similarly, if the limit is positive they would be all past incomplete, thus the limit is zero and so property (a) of Theorem 3.1 applies. ∎
The proof of the previous theorem uses results on the relationship between the behavior of the integral $\int\omega$ over the generators and the affine completeness of the generators, previously obtained in [19]. We stress that, as a consequence, the proof of the constancy of surface gravity in the degenerate case will not be independent of results obtained in the non-degenerate case. To the contrary, both the dichotomy and Theorem 3.3 depend on recent findings that originated in the study of the non-degenerate case.
@@ 4 Hodge theory and the isometric case
The exterior algebra $\Lambda^{*}T^{*}H$ is embedded in the covariant tensor algebra via antisymmetrization in such a way that, for 1-forms $\alpha,\beta$, the wedge product satisfies [37, 55, 32] (Spivak’s convention) $\alpha\wedge\beta=\alpha\otimes\beta-\beta\otimes\alpha$. This extends to arbitrary $k$-forms through multilinear antisymmetric combinations: For 1-forms $\alpha_{1},\dots,\alpha_{k}$,
@@
with $\delta\equiv 0$ on $\Omega^{0}(H)$. It relates to ${\rm d}$ and $\star$ via [48, Lemma 27]:
@@
A $k$-form $\omega$ is harmonic if $\Delta\omega=0$. On a compact manifold $H$, this condition simplifies significantly [48, Lemma 28][27]:
Theorem 4.1.
@@
Since (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is positive definite, $\Delta\omega=0$ implies ${\rm d}\omega=0$ and $\delta\omega=0$. The converse is immediate. ∎
@@
Theorem 4.2 (Hodge Decomposition).
@@
The following fact is well known, e.g. [14, Lemma 3.8]. We include the proof for completeness
Lemma 4.3.
@@
Proof.
Indeed, $L_{K}\alpha$ is closed as ${\rm d}L_{K}\alpha=L_{K}{\rm d}\alpha=0$ (here we did not use the Killing property), and also coclosed because $L_{K}$ commutes with $\delta$ (as $\delta$ depends just on the metric) and so $\delta L_{K}\alpha=L_{K}\delta\alpha=0$. As $L_{K}\alpha$ is both closed and coclosed, it is harmonic. But by Cartan’s formula $L_{K}\alpha={\rm d}(i_{K}\alpha)$ which proves that $L_{K}\alpha$ is both harmonic and exact, thus, by Hodge decomposition, $L_{K}\alpha=0$. ∎
@@
where $\alpha$ is a harmonic 1-form ($\Delta\alpha=0)$; $\beta$ is a co-exact, $\beta=\delta\eta$ for some 2-form $\eta$, and $f$ is a function on $H$. Since $\alpha$ is harmonic and $n$ is Killing we have by Lemma 4.3, Cartan’s magic formula and Thm. 4.1,
@@
which implies $\alpha(n)=a=\mbox{const}$. It is well-known that as $n$ is Killing, $L_{n}$ commutes with $\delta$ (as we recalled in the proof of Lemma 4.3), so $L_{n}\beta=L_{n}\delta\eta=\delta L_{n}\eta$ which shows that $L_{n}\beta$ is co-exact. From (23), we have ${\rm d}\omega={\rm d}\beta$ since ${\rm d}\alpha=0$ (a harmonic form is both closed and co-closed). Then we get $i_{n}{\rm d}\beta=i_{n}{\rm d}\omega=0$. Using Cartan’s magic formula, it follows that $L_{n}\beta={\rm d}(i_{n}(\beta))={\rm d}(\beta(n))$. We have proved that $L_{n}\beta$ is both co-exact and exact, which by Hodge $L_{2}$-orthogonal decomposition implies that $L_{n}\beta$ vanishes. So, we proved that $0=L_{n}\beta={\rm d}(\beta(n))$, which means that $\beta(n)=b=\mbox{const}$. Evaluating (23) over $n$ we get $\kappa=c+n(f)$ where $c:=a+b$ is a constant. ∎
Proposition 4.5.
If for a given spacetime dimension and under the dominant energy condition orientable degenerate compact horizons admit zero surface gravity, then the same is true in the non-orientable case.
If for a given spacetime dimension and under the dominant energy condition orientable degenerate compact horizons with isometric flow admit zero surface gravity, then the same is true in the non-orientable case.
Proof.
@@
On each element of the covering we have $\nabla n=n_{1}\otimes\omega_{1}+n_{2}\otimes\omega_{2}$ but also $\nabla n=n\otimes\omega$. Taking the trace of the former equation $\textrm{tr}(\nabla n)=\omega_{1}(n_{1})+\omega_{2}(n_{2})=0$ where we used vanishing of surface gravity on $\tilde{H}$, thus $0=\textrm{tr}(\nabla n)=\omega(n)$, which means vanishing of surface gravity with the choice $n=n_{1}+n_{2}$ all over $H$. ∎
@@
Theorem 4.6.
On a smooth spacetime $(M,g)$ suppose the dominant energy condition holds and let $H$ be a smooth compact degenerate horizon whose Riemannian flow is isometric. Then it admits a smooth tangent vector field of zero surface gravity (i.e. geodesic).
Proof.
By Prop. 4.5 it is sufficient to prove the statement for $H$ orientable. Since the flow is isometric we can choose a lightlike vector field $n$ which is Killing for some Riemannian metric $h$ on $H$. Let $\omega$ be the 1-form defined by Eq. (1). From the dominant energy condition and Prop. 2.2 we have $i_{n}{\rm d}\omega=0$. As all the assumptions of Theorem 4.4 are satisfied there are a smooth function $f$ and a constant $c$ such that $n(f)=\omega(n)-c=\kappa-c$. Setting $n^{\prime}=e^{-f}n$ we get from Eq. (3) that $\kappa^{\prime}=ce^{-f}$ which has the sign of $c$.
@@
Theorem 4.7.
On a smooth spacetime $(M,g)$ of dimension 2,3 or 4, suppose the dominant energy condition holds and let $H$ be a smooth compact degenerate horizon. Then it admits a smooth lightlike tangent vector field $n$ of zero surface gravity (i.e. geodesic).
Proof.
By Prop. 4.5 it is sufficient to consider the orientable case, and by Thm. 4.6 it is sufficient to consider the non-isometric case. In dimension 2 and 3 this concludes the proof as there are no orientable compact horizons with non-isometric Riemannian flow, see Thm. 1.5. In dimension 4 we are left with the structure of case (iii-b) of Theorem 1.4. This is explored in Carrière original paper [9] and [23, Example 2]. It is known as the hyperbolic torus $T^{3}_{A}$, where $A\in SL(2,\mathbb{Z})$, $\textrm{tr}A>2$, because it can be obtained as a suspension of base $T^{2}$ with the resulting flow being Anosov’s (roughly “suspension” here refers to the construction which goes as follows: consider the product $T^{2}\times[0,1]$ and identify the two $T^{2}$ bases after acting with $A$ on one of them). However, in Carrière non-isometric construction the Riemannian flow is not the natural one of the suspension but that of one of the two eigenvectors of $A$.
We continue with two independent arguments. The former uses the possibility of solving the cohomological equation via Fourier analysis over torus bundles, under a Diophantine condition, as shown by Dehghan-Nezhad and El Kacimi Alaoui [11]. The latter makes use of some deep and powerful results by Álvarez López and Domínguez on the possibility of finding a bundle like metric with a basic mean curvature form $\mu$ (we recalled this concept in our previous work [23]). In the next section we shall really obtain a third proof as the 4-dimensional case will follow from the general case.
@@
    We have shown in [23, Example 2] in some detail, that the two roots $\lambda_{+},\lambda_{-}$ of the equation $\lambda^{2}-\textrm{tr}(A)\lambda+1$ are irrational. We set $v_{+}=(a,b)^{T}$ for one eigenvector of $A$ and by imposing $A(v_{+})=\lambda_{+}v_{+}$, we find $a,b\neq 0$ with an irrational ratio $c=b/a$. A similar conclusion is reached considering $\lambda_{-}$ and $v_{-}$. As any eigenvector can be rescaled, we can consider $v_{+}=(1,c)$ where the first line of $A(v_{+})=\lambda_{+}v_{+}$ gives $A_{11}+cA_{12}=\lambda_{+}$. But $A$ has integer coefficients so $A_{12}\neq 0$ and hence $c=\frac{\lambda_{+}-A_{11}}{A_{12}}$. As all the numbers on the right-hand side are algebraic so is $c$. Roth’s theorem [52, 22] states that for every irrational algebraic number $c$ there are positive constants $\tau>1$ and $C<1$ such that for every integers $p,q$, $q\neq 0$, $|c+p/q|\geq C/|q|^{1+\tau}$. As a consequence, for every irrational algebraic number $c$ there are positive constants $\tau>1$ and $C<1$ such that for every integers $p,q$, $|p+qc|\geq C/\max(|p|,|q|)^{\tau}$, which implies that for every vector $z\in\mathbb{Z}^{2}$, $|z\cdot v_{+}|\geq C/\max(|z_{1}|,|z_{2}|)^{\tau}$. In other words, $v_{+}$ is a Diophantine vector [54, 11]. An analogous conclusion holds for $v_{-}$.
    By Remark 2.1 it is sufficient to prove that the cohomological equation can be solved over $T^{3}_{A}$ endowed with Carrière’s non-isometric flow. Let $n$ be the vector field generating the flow as provided in coordinates, for instance, in [23, Ex. 2] (there it is denoted $Y$ and $u$ is denoted $x$). There it is shown that in the local trivialization it is the constant vector $v_{\pm}$ up to a factor dependent on $u$.
@@
    This geometry, in which we have a torus bundle over a manifold (connected and orientable but not necessarily compact), and a Riemannian flow which can be represented, in a local trivialization of the bundle, by a Diophantine linear flow on the torus fiber independent of the coordinate(s) $u$ of the base, has been considered by Dehghan-Nezhad and El Kacimi Alaoui [11]. Their result [11, Thm. 2.4, also p. 1114-15] states that if $n^{\prime}$ is a Diophantine vector field generating the Riemannian flow, then the cohomological equation $n^{\prime}(f)=\kappa^{\prime}$ has a smooth solution for any smooth $\kappa^{\prime}$ such that the integral of $\kappa^{\prime}$ on every torus fiber vanishes. Here the measure to be used is the canonical one of the torus in the trivialization. As explained in the previous paragraphs, we just need to apply their theorem to $I\times T^{2}$, and the same paragraphs show that the integral assumption is satisfied. By multiplying $n^{\prime}(f)=\kappa^{\prime}$ by $r(u)$ we get that the solution $f$ also solves $n(f)=\kappa$, and now we can use the partition of unity to globalize.
    In conclusion, the cohomological equation $n(f)=\kappa$ over $T^{3}_{A}$ can be solved for $\kappa=\omega(n)$, and so the horizon admits a smooth lightlike tangent field of zero surface gravity cf. Remark 2.1.

### v2

@@ 2 The cohomological equation
The foliated cohomology can be defined for more general foliated manifolds but is particularly simple to describe in this flow setting dehghan07 . Let $\chi$ be a 1-form whose kernel is supplementary at each point to the tangent space $D:=\textrm{span}(n)$ of the leaves and $\chi(n)=1$. Then
@@
The literature on the cohomological equation is extensive but results are focused on proving the existence of $f$ for any $\kappa$ up to some normalization conditions to be imposed on the integral of $\kappa$. In general, the problem can be solved under restrictive conditions on the flow, in fact Katok’s conjecture (now proved for $H$ of dimension 3 matsumoto09 ; kocsard09 ; forni08 , which is the physical 4-dimensional spacetime case) states that on a connected orientable compact manifold $H$ the equation $n(f)=\kappa-c(\kappa)$ can be solved for any function $\kappa$ and for some constant $c(\kappa)$, only if the flow is conjugate to a Diophantine linear flow on the torus.
Instead, in our spacetime problem we need to solve the cohomological equation for just one function $\kappa$. We do not expect severe constraints on the topology of $H$ or on the flow but we have two simplifications on our side. On the one hand, we already know the flow to be Riemannian and, on the other hand, the function $\kappa$ is not any function as it reads $\kappa=\omega(n)$ i.e. it is derived from a 1-form $\omega$ with special properties. We shall need to use these facts. In particular, we shall need that under the null energy condition, on the horizon $H$, ${\rm d}\omega(n,\cdot)=0$ (Cor. 1)
As we shall see, previous studies on the cohomological equation dehghan07 will really be used just in dealing with the exceptional non-isometric case $T^{3}_{A}$ (see case (iii-b) of Thm. 1.2) as it involves a Diophantine linear flow on the torus.
@@ 3 Consequences of ergodic theory and ribbon argument
Proof
The flow $(\phi_{t})_{t\in\mathbb{R}}$ of $n$ is a measure preserving flow since it preserves the Riemannian volume element, in fact, as shown in the proof of (minguzzi24c, , Prop. 22) $\widetilde{\textrm{div}}(n)=0$. Let us denote $\kappa=\omega(n)$. Then, Birkhoff’s Ergodic Theorem for measure preserving transformations gives birkhoff31 ; nemytskii60 ; arnold68
@@
It can be observed that the argument of the limit is bounded as $\omega(n)$ is a continuous function on $H$, hence bounded. The important point is that by Birkhoff’s theorem this limit converges for some point. Let $p$ be one such point and suppose that the limit is zero. Let $n^{*}$ be a smooth 1-form field such that $n^{*}(n)=1$. A curve with tangent in $\ker n^{*}$ is said to be $n^{*}$-horizontal. The limit $\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\omega(\dot{x}){\rm d}t$ does not really depend on the starting point of the curve, so for any sufficiently close integral curve $x^{\prime}$ we can redefine the starting point $x^{\prime}(0)$ of the parametrization so that there is a horizontal curve $\sigma_{0}$ connecting $x(0)$ to $x^{\prime}(0)$ (this can be done so that its projection on a local cylinder neighborhood is a geodesic for $g_{T}$, see minguzzi21 ). Now, it is possible to construct a ribbon from the curves $x|_{[0,t]}$, $\sigma_{0}|_{[0,1]}$ and $x^{\prime}|_{[0,t^{\prime}(t)]}$, and application of Stokes’s theorem and the bound described in minguzzi21 implies that
@@
The bound on the right-hand side does not depend on how elongated is the ribbon (see (minguzzi21, , Eq. (18)-(19)) and take into account that this conclusion only depends on $L_{n}g_{T}=0$ not on the spacetime interpretation of the various objects in that work) and in (minguzzi21, , App.) it is detailed how $t\to+\infty$ implies $t^{\prime}\to\infty$, and conversely. In fact, on the same appendix (footnote 6) it is shown that there is a constant $C>0$ such that $te^{-C}\leq t^{\prime}\leq te^{C}$. Dividing the previous equation by $t$ and taking the limit we get, as $t^{\prime}/t$ remains bounded away from zero, $\lim_{t^{\prime}\to\infty}\frac{1}{t^{\prime}}\int_{x^{\prime}([0,t^{\prime}])}\omega=0$. This shows that the set $O$ of points $p$ for which the orbit starting from it has first integral existing and equal to zero is open. In fact it is also closed, for if $q\in\partial O$ then for every neighborhood $U\ni q$, we can find some $p\in O$ for which the integral exists and vanishes. Then, reapplying the ribbon argument, we get that the same is true for the orbit from $q$. By the connectedness of $H$ we conclude $O=H$. By Birkhoff’s theorem $\int_{H}\omega(n){\rm d}\tilde{g}=0$. This shows that (a) implies its strengthened form and (c).
@@
Let us prove (a) implies (d). In the proof of (minguzzi24c, , Prop. 24) we proved that for $N=H$ or $N$ equal to a torus closure of an orbit, the restricted flow of $n$ is volume preserving. Actually, the proof applies unaltered for any saturated submanifold $N$. The Birkhoff’s Ergodic Theorem can thus be applied to the submanifold which proves (d). Conversely, when the last equation of point (d) is applied to $N=H$ we get point (c), thus (d) implies (c).
Remark 3
@@
Theorem 3.2
Suppose that the spacetime satisfies the null energy condition and let $n$ be any smooth lightlike field tangent to a compact degenerate horizon $H$. Let $\omega$ be the 1-form defined by Eq. (1). Then interpreting $H$ as a Riemannian flow structure, the properties of Theorem 3.1 are satisfied.
Proof
The limit $\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\kappa(x(t)){\rm d}t$ is known to exist for some generator but it cannot be different from zero. Indeed, if it is negative then for every generator $\int_{0}^{t}\kappa(x(t)){\rm d}t=-\infty$ and by the equivalence proved in (minguzzi21, , Def. 4) the generators of the horizons would be all future incomplete. Similarly, if the limit is positive they would be all past incomplete, thus the limit is zero and so property (a) of Theorem 3.1 applies.
The proof of the previous theorem uses results on the relationship between the behavior of the integral $\int\omega$ over the generators and the affine completeness of the generators, previously obtained in minguzzi21 . We stress that, as a consequence, the proof of the constancy of surface gravity in the degenerate case will not be independent of results obtained in the non-degenerate case. To the contrary, both the dichotomy and Theorem 3.2 depend on recent findings that originated in the study of the non-degenerate case.
@@ 4 Hodge theory and the isometric case
The exterior algebra $\Lambda^{*}T^{*}H$ is embedded in the covariant tensor algebra via antisymmetrization in such a way that, for 1-forms $\alpha,\beta$, the wedge product satisfies misner73 ; spivak79 ; lee13 (Spivak’s convention) $\alpha\wedge\beta=\alpha\otimes\beta-\beta\otimes\alpha$. This extends to arbitrary $k$-forms through multilinear antisymmetric combinations: For 1-forms $\alpha_{1},\dots,\alpha_{k}$,
@@
with $\delta\equiv 0$ on $\Omega^{0}(H)$. It relates to ${\rm d}$ and $\star$ via (petersen06, , Lemma 27):
@@
A $k$-form $\omega$ is harmonic if $\Delta\omega=0$. On a compact manifold $H$, this condition simplifies significantly (petersen06, , Lemma 28)jost11 :
Theorem 4.1
@@
Since (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is positive definite, $\Delta\omega=0$ implies ${\rm d}\omega=0$ and $\delta\omega=0$. The converse is immediate.
@@
Theorem 4.2 (Hodge Decomposition)
@@
The following fact is well known, e.g. (egidi23, , Lemma 3.8). We include the proof for completeness
Lemma 1
@@
Proof
Indeed, $L_{K}\alpha$ is closed as ${\rm d}L_{K}\alpha=L_{K}{\rm d}\alpha=0$ (here we did not use the Killing property), and also coclosed because $L_{K}$ commutes with $\delta$ (as $\delta$ depends just on the metric) and so $\delta L_{K}\alpha=L_{K}\delta\alpha=0$. As $L_{K}\alpha$ is both closed and coclosed, it is harmonic. But by Cartan’s formula $L_{K}\alpha={\rm d}(i_{K}\alpha)$ which proves that $L_{K}\alpha$ is both harmonic and exact, thus, by Hodge decomposition, $L_{K}\alpha=0$.
@@
where $\alpha$ is a harmonic 1-form ($\Delta\alpha=0)$; $\beta$ is a co-exact, $\beta=\delta\eta$ for some 2-form $\eta$, and $f$ is a function on $H$. Since $\alpha$ is harmonic and $n$ is Killing we have by Lemma 1, Cartan’s magic formula and Thm. 4.1,
@@
which implies $\alpha(n)=a=\mbox{const}$. It is well-known that as $n$ is Killing, $L_{n}$ commutes with $\delta$ (as we recalled in the proof of Lemma 1), so $L_{n}\beta=L_{n}\delta\eta=\delta L_{n}\eta$ which shows that $L_{n}\beta$ is co-exact. From (24), we have ${\rm d}\omega={\rm d}\beta$ since ${\rm d}\alpha=0$ (a harmonic form is both closed and co-closed). Then we get $i_{n}{\rm d}\beta=i_{n}{\rm d}\omega=0$. Using Cartan’s magic formula, it follows that $L_{n}\beta={\rm d}(i_{n}(\beta))={\rm d}(\beta(n))$. We have proved that $L_{n}\beta$ is both co-exact and exact, which by Hodge $L_{2}$-orthogonal decomposition implies that $L_{n}\beta$ vanishes. So, we proved that $0=L_{n}\beta={\rm d}(\beta(n))$, which means that $\beta(n)=b=\mbox{const}$. Evaluating (24) over $n$ we get $\kappa=c+n(f)$ where $c:=a+b$ is a constant.
Proposition 3
If for a given spacetime dimension and under the null energy condition orientable degenerate compact horizons admit zero surface gravity, then the same is true in the non-orientable case.
If for a given spacetime dimension and under the null energy condition orientable degenerate compact horizons with isometric flow admit zero surface gravity, then the same is true in the non-orientable case.
Proof
@@
On each element of the covering we have $\nabla n=n_{1}\otimes\omega_{1}+n_{2}\otimes\omega_{2}$ but also $\nabla n=n\otimes\omega$. Taking the trace of the former equation $\textrm{tr}(\nabla n)=\omega_{1}(n_{1})+\omega_{2}(n_{2})=0$ where we used vanishing of surface gravity on $\tilde{H}$, thus $0=\textrm{tr}(\nabla n)=\omega(n)$, which means vanishing of surface gravity with the choice $n=n_{1}+n_{2}$ all over $H$.
@@
Theorem 4.4
On a smooth spacetime $(M,g)$ suppose the null energy condition holds and let $H$ be a smooth compact degenerate horizon whose Riemannian flow is isometric. Then it admits a smooth tangent vector field of zero surface gravity (i.e. geodesic).
Proof
By Prop. 3 it is sufficient to prove the statement for $H$ orientable. Since the flow is isometric we can choose a lightlike vector field $n$ which is Killing for some Riemannian metric $h$ on $H$. Let $\omega$ be the 1-form defined by Eq. (1). From the null energy condition and Cor. 1 we have $i_{n}{\rm d}\omega=0$. As all the assumptions of Theorem 4.3 are satisfied there are a smooth function $f$ and a constant $c$ such that $n(f)=\omega(n)-c=\kappa-c$. Setting $n^{\prime}=e^{-f}n$ we get from Eq. (3) that $\kappa^{\prime}=ce^{-f}$ which has the sign of $c$.
@@
Theorem 4.5
On a smooth spacetime $(M,g)$ of dimension 2,3 or 4, suppose the null energy condition holds and let $H$ be a smooth compact degenerate horizon. Then it admits a smooth lightlike tangent vector field $n$ of zero surface gravity (i.e. geodesic).
Proof
By Prop. 3 it is sufficient to consider the orientable case, and by Thm. 4.4 it is sufficient to consider the non-isometric case. In dimension 2 and 3 this concludes the proof as there are no orientable compact horizons with non-isometric Riemannian flow, see Thm. 1.3. In dimension 4 we are left with the structure of case (iii-b) of Theorem 1.2. This is explored in Carrière original paper carriere84 and (minguzzi24c, , Example 2). It is known as the hyperbolic torus $T^{3}_{A}$, where $A\in SL(2,\mathbb{Z})$, $\textrm{tr}A>2$, because it can be obtained as a suspension of base $T^{2}$ with the resulting flow being Anosov’s (roughly “suspension” here refers to the construction which goes as follows: consider the product $T^{2}\times[0,1]$ and identify the two $T^{2}$ bases after acting with $A$ on one of them). However, in Carrière non-isometric construction the Riemannian flow is not the natural one of the suspension but that of one of the two eigenvectors of $A$.
We continue with two independent arguments. The former uses the possibility of solving the cohomological equation via Fourier analysis over torus bundles, under a Diophantine condition, as shown by Dehghan-Nezhad and El Kacimi Alaoui dehghan07 . The latter makes use of some deep and powerful results by Álvarez López and Domínguez on the possibility of finding a bundle like metric with a basic mean curvature form $\mu$ (we recalled this concept in our previous work minguzzi24c ). In the next section we shall really obtain a third proof as the 4-dimensional case will follow from the general case.
@@
    We have shown in (minguzzi24c, , Example 2) in some detail, that the two roots $\lambda_{+},\lambda_{-}$ of the equation $\lambda^{2}-\textrm{tr}(A)\lambda+1$ are irrational. We set $v_{+}=(a,b)^{T}$ for one eigenvector of $A$ and by imposing $A(v_{+})=\lambda_{+}v_{+}$, we find $a,b\neq 0$ with an irrational ratio $c=b/a$. A similar conclusion is reached considering $\lambda_{-}$ and $v_{-}$. As any eigenvector can be rescaled, we can consider $v_{+}=(1,c)$ where the first line of $A(v_{+})=\lambda_{+}v_{+}$ gives $A_{11}+cA_{12}=\lambda_{+}$. But $A$ has integer coefficients so $A_{12}\neq 0$ and hence $c=\frac{\lambda_{+}-A_{11}}{A_{12}}$. As all the numbers on the right-hand side are algebraic so is $c$. Roth’s theorem roth55 ; hindry00 states that for every irrational algebraic number $c$ there are positive constants $\tau>1$ and $C<1$ such that for every integers $p,q$, $q\neq 0$, $|c+p/q|\geq C/|q|^{1+\tau}$. As a consequence, for every irrational algebraic number $c$ there are positive constants $\tau>1$ and $C<1$ such that for every integers $p,q$, $|p+qc|\geq C/\max(|p|,|q|)^{\tau}$, which implies that for every vector $z\in\mathbb{Z}^{2}$, $|z\cdot v_{+}|\geq C/\max(|z_{1}|,|z_{2}|)^{\tau}$. In other words, $v_{+}$ is a Diophantine vector schmidt80 ; dehghan07 . An analogous conclusion holds for $v_{-}$.
    By Remark 2 it is sufficient to prove that the cohomological equation can be solved over $T^{3}_{A}$ endowed with Carrière’s non-isometric flow. Let $n$ be the vector field generating the flow as provided in coordinates, for instance, in (minguzzi24c, , Ex. 2) (there it is denoted $Y$ and $u$ is denoted $x$). There it is shown that in the local trivialization it is the constant vector $v_{\pm}$ up to a factor dependent on $u$.
@@
    This geometry, in which we have a torus bundle over a manifold (connected and orientable but not necessarily compact), and a Riemannian flow which can be represented, in a local trivialization of the bundle, by a Diophantine linear flow on the torus fiber independent of the coordinate(s) $u$ of the base, has been considered by Dehghan-Nezhad and El Kacimi Alaoui dehghan07 . Their result (dehghan07, , Thm. 2.4, also p. 1114-15) states that if $n^{\prime}$ is a Diophantine vector field generating the Riemannian flow, then the cohomological equation $n^{\prime}(f)=\kappa^{\prime}$ has a smooth solution for any smooth $\kappa^{\prime}$ such that the integral of $\kappa^{\prime}$ on every torus fiber vanishes. Here the measure to be used is the canonical one of the torus in the trivialization. As explained in the previous paragraphs, we just need to apply their theorem to $I\times T^{2}$, and the same paragraphs show that the integral assumption is satisfied. By multiplying $n^{\prime}(f)=\kappa^{\prime}$ by $r(u)$ we get that the solution $f$ also solves $n(f)=\kappa$, and now we can use the partition of unity to globalize.
    In conclusion, the cohomological equation $n(f)=\kappa$ over $T^{3}_{A}$ can be solved for $\kappa=\omega(n)$, and so the horizon admits a smooth lightlike tangent field of zero surface gravity cf. Remark 2.